(please give brief explanation to this sentence -> two squares have sides p cm and (p+5) cms)
two squares have sides p cm and (p+5) cms. the sum of theirs squares is 625 sq.cm. the sides of the squares are

Answer by ankor@dixie-net.com(22740)

two squares have sides p cm and (p+5) cms. the sum of theirs squares is 625 sq.cm. the sides of the squares are
:
p^2 + (p+5)^2 = 625
FOIL (p+5)(p+5)
p^2 + p^2 + 5p + 5p + 25 - 625 = 0
combine like terms
2p^2 + 10p - 600 = 0
Simplify divide by 2
p^2 + 5p - 300 = 0
Factors to
(p+20)(p-15) = 0
The positive solution is all we want here
p = 15 cm the side of smaller square
:
you find the length of the larger square
:
Then you can check this: 15^2 + 20^2 = ?